OpenStudy (anonymous):

Find an equation of the tangent line to y=f(x) at x=1 f(x)=3e^(x^2) please help1!!

OpenStudy (anonymous):

Have you found the derivative of the curve yet?

OpenStudy (solomonzelman):

1) Derive the equation. 2) Plug in the 1 for x, (since it is at x=1) 3) then the point is (1,f(1)).

OpenStudy (solomonzelman):

Any questions ?

OpenStudy (solomonzelman):

I mean plug in 1 for x, into the derivative.

OpenStudy (solomonzelman):

Step 1: \(\large\color{blue}{ y=3e^{x^2} }\) \(\large\color{blue}{ y'=3e^{x^2}\times2x=6xe^{x^2} }\)

OpenStudy (solomonzelman):

Step 2: \(\large\color{blue}{ y'=3e^{x^2}\times2x=6xe^{x^2} }\) \(\large\color{blue}{ m_{tan}=6(1)e^{1^2}=6e^1=6e }\)

OpenStudy (solomonzelman):

So you know the slope is 6e.

OpenStudy (solomonzelman):

Find f(1), which is the y-coordinate of the point.

OpenStudy (solomonzelman):

\(\large\color{blue}{ f(x)=3e^{x^2}}\) \(\large\color{blue}{ f(1)=3e^{1^2}=3e}\) So the point is (1,3e)

OpenStudy (solomonzelman):

\(\large\color{blue}{ y-y_1=m(x-x_1)}\) \(\large\color{blue}{ y-3e=6e(x-1)}\) \(\large\color{blue}{ y-3e=6e~x~-~6e}\) \(\large\color{blue}{ y=6e~x~-~3e}\)

OpenStudy (solomonzelman):

This is the answer. ( \(\large\color{blue}{ y=6e~x~-~3e}\) )

OpenStudy (anonymous):

Oh my, thank you sosososososo mcuh!!!!!!!!! :D